OpenStudy (anonymous):

If tan theta = -5/7, cos theta <0, what is sin, cos, csc,and cot theta

OpenStudy (anonymous):

tan inverse(-5/7)=-35.537 so angle is -35.537

OpenStudy (anonymous):

h^2=root(49+25)=root(74) =root(74) sin(theta)=-5/(root 74) cos(theta)=7/(root 74)

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

how does one solve this partial-fraction decomposition : 3x^3-5x^2 +12 +4/ x^4 -16

OpenStudy (anonymous):

partial-fraction decomposition is related to complex numbers process called residues if you are familiar with it you should be okey

OpenStudy (anonymous):

No am not familiar with residues

OpenStudy (anonymous):

ok so \[3x^3-5x^2+12x+4/(x^2-4)(x^2+4)\] \[form: AX+B/(x^2-4) +CX+D/(x^2+4)\] then multiplying through LHS with (x^2+4)(x^2-4)

OpenStudy (anonymous):

equate the like terms and get value of A,B,C,D